40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15157     Entry time: Sun Jan 26 14:40:55 2020
Author: gautam 
Type: Update 
Category: ALS 
Subject: ALS OOL noise 

In preparation for resuming IFO locking activities, I measured the ALS noise with the arm lengths locked to IR, AUX laser frequencies locked to the arm lengths. Looks promising (y-axis units are Hz/rtHz).

Attachment 1: ALSnoise_20200126.pdf  77 kB  | Hide | Hide all
ALSnoise_20200126.pdf
ELOG V3.1.3-